OpenStudy (anonymous):

Find the limit as x approaches 0 for (t^2)/(〖1-cos〗^2 t)

OpenStudy (campbell_st):

I think it should be \[\lim_{t \rightarrow 0} \frac{t^2}{(1 - \cos(?))^2 t}\] but if its as above you can remove a common factor of t from the numerator and denominator.

OpenStudy (anonymous):

its not that that one haha its \[\frac{ t ^{2} }{ 1-\cos ^{2}t }\]

OpenStudy (anonymous):

\[\huge \frac{t^2}{\sin^2t}=\frac{1}{(\sin^2t )\times \frac{1}{t^2}}=\frac{1}{\frac{\sin^2t}{t^2}}\] \[\lim_{t \rightarrow 0} \frac{\sin t}{t}=0\]

OpenStudy (anonymous):

the answer is suppose to be 1...

OpenStudy (anonymous):

sry \[\frac{\sin t}{t}=1\] \[\frac{1}{1^2}=1\] yes

OpenStudy (anonymous):

oh i see it, thanks!